# Era Splitting - Invariant Learning for Decision Trees

### v4

@@ 5 Experimental Methods
Four experiments have been designed to empirically validate the newly proposed splitting criteria. The first experiment is an original design, called the shifted sine wave, validating the era splitting models in a simple one-dimensional setting. The second and third experiments are binary classification problems from the OOD literature: the synthetic memorization data set presented in ([37]) and the Camelyon17 (CAncer MEtastases in LYmph nOdes challeNge) data set from ([23]). These two experiments confirm the new splitting criteria are effective in known OOD settings. The forth is the target application of this research, the Numerai data set.
The experiments follow a similar procedure for each. Random configurations are drawn from a grid of parameter values. The parameters include the number of boosting iterations, maximum number of leaves, learning rate, etc. All relevant parameters in Scikit-Learn’s HistGradientBoostingRegressor ([38]) (the baseline model) are available. The number of random configurations (between 15 and 30) and the exact parameter ranges vary slightly from one data set to another, depending on training time and specifics of each data set. For the full list of parameters and example values, see table 2. See the code for details, referenced at the end of this subsection. Each of the three experimental models are paired with each random configuration: original, era splitting, and directional era splitting. In this way all three models are trained and evaluated over the same set of parameter values. In each case, the model is trained on the the training set and evaluated on the OOS test set. Evaluation statistics are recorded for each configuration. Accuracy is the key statistic for classification problems. The regressor is converted to a binary classifier by rounding the predictions to the nearest integer (0 or 1). For regression problems the mean-squared error (MSE) and Pearson correlation (Corr) are observed.
All the code for reproducing the experiments presented in this paper is available in the following two links. The first contains the forked version of Scikit-Learn[38] which implements the era splitting models, and the second contains all the notebooks to recreate the figures and experiments from this research.
@@ 5.2 Synthetic Memorization Data Set
The synthetic memorization data set was developed in ([37]) specifically to investigate an ML model’s ability to learn invariant predictors in high-dimensional data in the presence of confounding spurious signals. It is a supervised binary classification task with target labels taking the values $\{0,1\}$. The data set is composed such that the first two input dimensions encode the invariant signal, which is an interlaced spiral shape, where one of the arms belongs to one class, and the other arm the other class. The remaining input dimensions are used to encode simple shortcuts, or spurious easy-to-learn signals. These are simply two clusters of data points in high dimensions, where one cluster belongs to one class and the other the other class. The clusters are linearly separable, being easy to learn. The spurious aspect is realized through a mechanism where each era (environment) of data has a shift in the location of the two clusters. The shift happens in such a way that, when all the environments of the training data set are pooled together, the shortcut is still present. Refer to figure 5 or figure 5 from ([37]) for a visualization of the data in four dimensions. When environments 1 & 2 are pooled together, there is still an easily identifiable linear decision boundary present in the spurious signal feature dimensions.
Figure 5: The synthetic memorization data set from ([37]). Each environment contains a random sample of the invariant spiral signal along the first two dimensions. The third and forth dimension contain a simple linearly-separable spurious signal that changes with each environment. When pooling environments together, the pooled spurious signals still create a simple linear decision boundary. During OOS testing, the spurious signal is replaced by random noise.
During test time the spurious signal is replaced by random noise. Any model that bases its predictions solely on the spurious signal will be worthless when evaluated on the test data. If a model has actually learned the invariant, albeit harder-to-learn signal, it will still perform well on the test set. The authors of ([37]) developed a technique called the AND-mask, which is an alteration to the naive gradient descent algorithm, incorporating the era-wise information. They were able to show that a naive MLP model simply learns the spurious signals in the data, and performs poorly on the test data. Meanwhile their AND-mask model was able to learn in such a way where it still performs well (98% accuracy) on the test data.
@@ 5.3 Camelyon17 Data Set
The Camelyon17 data set ([23]) is an example of a real-world dataset exhibiting distributional shifts. This is one of the Wilds data sets that is distributed and supported by a team at Stanford ([33]). This data set aims at the automation of breast cancer detection in medical settings. The input data is a $96\times 96$ pixel histopathological image and the target label is a binary indicator that is positive if the central $32\times 32$ pixel region contains any tumor tissue. There are variations in data generation and processing technique from one hospital to another. The domain generalization problem is to learn to detect tumors from images from a number of hospitals, and have this detection ability transfer to new hospitals not in the training set.
@@ 6 Results
The results for all four experiments are summarized in table 3 and figure 6. Table 3 tabulates the best evaluation score over all the random configurations for each model split type. In each experiment the best model is either an era splitting model or a directional era splitting model. Figure 6 displays box plots of the evaluation metrics for both training and test sets over all the model configurations, separated by split type. A consistent trend emerges. The original split models are characterized by the best in-sample performance and the worst OOS performance, ie. the largest generalization gap. The era split and directional era split models have decreased in-sample and increased OOS performance, a smaller generalization gap, consistently in every experiment. This supports the claim of proposition 4.3, that these new splitting criteria act as effective regularizers.
@@
Figures 7 and 8 visualize the decision surfaces as a function of input data values for the shifted sine wave and synthetic memorization data sets. The sine wave experiment is one dimensional and easy to plot. The original split criterion results in a much noisier decision boundary than both era splitting and directional era splitting for this configuration. The synthetic memorization data set is high-dimensional, with the invariant swirl signal encoded on the first two dimensions. Plotting the predicted values as a function of these first two dimensions reveals whether or not the model has learned to use this signal. The original split model has been unable to learn the swirl signal, as well the predicted values are very extreme, meaning the model is very confident in those predictions. The high in-sample and low OOS accuracy scores indicate that these models have only learned the spurious signals which disappear in the OOS test set. On the other hand, both the experimental models have clearly learned the swirl signal, with higher predicted values along one arm and lower values along the other.
@@
All the results have been check-pointed in iPython notebooks available at the example code repository for this project:
github.com/jefferythewind/era-splitting-notebook-examples.
@@ 7 Discussion
One of the biggest obstacles to the wide-spread adoption of this model is the added time complexity that is added by the era splitting routine. Indeed, one of the main issues with GBDTs in general is that, in order to grow a tree, each node must compute the split score (equation 4) for every potential split in the entire data set. The calculation must be made for every distinct value for every feature in the data. The histogram method ([26]) in modern software libraries has helped by reducing the number of unique values in the features down to a set parameter value. Setting this parameter to a low value, like 5, can greatly reduce training time. The colsample bytree method also reduces the total number of computations needed by randomly sampling a subset of the features to use to build each tree, similar to the way bagging can reduce the number of rows.

### v5

@@ 5 Experimental Methods
Four experiments have been designed to empirically validate the newly proposed splitting criteria. The first experiment is an original design, called the shifted sine wave, validating the era splitting models in a simple one-dimensional setting. The second and third experiments are binary classification problems from the OOD literature: the synthetic memorization data set presented in Parascandolo et al. [2020] and the Camelyon17 (CAncer MEtastases in LYmph nOdes challeNge) data set from Bandi et al. [2018]. These two experiments confirm the new splitting criteria are effective in known OOD settings. The forth is the target application of this research, the Numerai data set.
The experiments follow a similar procedure for each. Random configurations are drawn from a grid of parameter values. The parameters include the number of boosting iterations, maximum number of leaves, learning rate, etc. All relevant parameters in Scikit-Learn’s HistGradientBoostingRegressor Pedregosa et al. [2011] (the baseline model) are available. The number of random configurations (between 15 and 30) and the exact parameter ranges vary slightly from one data set to another, depending on training time and specifics of each data set. For the full list of parameters and example values, see table 2. See the code for details, referenced at the end of this subsection. Each of the three experimental models are paired with each random configuration: original, era splitting, and directional era splitting. In this way all three models are trained and evaluated over the same set of parameter values. In each case, the model is trained on the the training set and evaluated on the OOS test set. Evaluation statistics are recorded for each configuration. Accuracy is the key statistic for classification problems. The regressor is converted to a binary classifier by rounding the predictions to the nearest integer (0 or 1). For regression problems the mean-squared error (MSE) and Pearson correlation (Corr) are observed.
@@ 5.2 Synthetic Memorization Data Set
The synthetic memorization data set was developed in Parascandolo et al. [2020] specifically to investigate an ML model’s ability to learn invariant predictors in high-dimensional data in the presence of confounding spurious signals. It is a supervised binary classification task with target labels taking the values $\{0,1\}$. The data set is composed such that the first two input dimensions encode the invariant signal, which is an interlaced spiral shape, where one of the arms belongs to one class, and the other arm the other class. The remaining input dimensions are used to encode simple shortcuts, or spurious easy-to-learn signals. These are simply two clusters of data points in high dimensions, where one cluster belongs to one class and the other the other class. The clusters are linearly separable, being easy to learn. The spurious aspect is realized through a mechanism where each era (environment) of data has a shift in the location of the two clusters. The shift happens in such a way that, when all the environments of the training data set are pooled together, the shortcut is still present. Refer to figure 6 or figure 5 from Parascandolo et al. [2020] for a visualization of the data in four dimensions. When environments 1 & 2 are pooled together, there is still an easily identifiable linear decision boundary present in the spurious signal feature dimensions.
Figure 6: The synthetic memorization data set from Parascandolo et al. [2020]. Each environment contains a random sample of the invariant spiral signal along the first two dimensions. The third and forth dimension contain a simple linearly-separable spurious signal that changes with each environment. When pooling environments together, the pooled spurious signals still create a simple linear decision boundary. During OOS testing, the spurious signal is replaced by random noise.
During test time the spurious signal is replaced by random noise. Any model that bases its predictions solely on the spurious signal will be worthless when evaluated on the test data. If a model has actually learned the invariant, albeit harder-to-learn signal, it will still perform well on the test set. The authors of Parascandolo et al. [2020] developed a technique called the AND-mask, which is an alteration to the naive gradient descent algorithm, incorporating the era-wise information. They were able to show that a naive MLP model simply learns the spurious signals in the data, and performs poorly on the test data. Meanwhile their AND-mask model was able to learn in such a way where it still performs well (98% accuracy) on the test data.
@@ 5.3 Camelyon17 Data Set
The Camelyon17 data set Bandi et al. [2018] is an example of a real-world dataset exhibiting distributional shifts. This is one of the Wilds data sets that is distributed and supported by a team at Stanford Koh et al. [2021]. This data set aims at the automation of breast cancer detection in medical settings. The input data is a $96\times 96$ pixel histopathological image and the target label is a binary indicator that is positive if the central $32\times 32$ pixel region contains any tumor tissue. There are variations in data generation and processing technique from one hospital to another. The domain generalization problem is to learn to detect tumors from images from a number of hospitals, and have this detection ability transfer to new hospitals not in the training set.
@@ 6 Results
The results for all four experiments are summarized in table 3 and figure 7. Table 3 tabulates the best evaluation score over all the random configurations for each model split type. In each experiment the best model is either an era splitting model or a directional era splitting model. Figure 7 displays box plots of the evaluation metrics for both training and test sets over all the model configurations, separated by split type. A consistent trend emerges. The original split models are characterized by the best in-sample performance and the worst OOS performance, ie. the largest generalization gap. The era split and directional era split models have decreased in-sample and increased OOS performance, a smaller generalization gap, consistently in every experiment. This supports the claim of proposition 5, that these new splitting criteria act as effective regularizers.
@@
Figures 8 and 9 visualize the decision surfaces as a function of input data values for the shifted sine wave and synthetic memorization data sets. The sine wave experiment is one dimensional and easy to plot. The original split criterion results in a much noisier decision boundary than both era splitting and directional era splitting for this configuration. The synthetic memorization data set is high-dimensional, with the invariant swirl signal encoded on the first two dimensions. Plotting the predicted values as a function of these first two dimensions reveals whether or not the model has learned to use this signal. The original split model has been unable to learn the swirl signal, as well the predicted values are very extreme, meaning the model is very confident in those predictions. The high in-sample and low OOS accuracy scores indicate that these models have only learned the spurious signals which disappear in the OOS test set. On the other hand, both the experimental models have clearly learned the swirl signal, with higher predicted values along one arm and lower values along the other.
@@ 7 Discussion
One of the biggest obstacles to the wide-spread adoption of this model is the added time complexity that is added by the era splitting routine. Indeed, one of the main issues with GBDTs in general is that, in order to grow a tree, each node must compute the split score (equation 26) for every potential split in the entire data set. The calculation must be made for every distinct value for every feature in the data. The histogram method Chen and Guestrin [2016a] in modern software libraries has helped by reducing the number of unique values in the features down to a set parameter value. Setting this parameter to a low value, like 5, can greatly reduce training time. The colsample bytree method also reduces the total number of computations needed by randomly sampling a subset of the features to use to build each tree, similar to the way bagging can reduce the number of rows.
@@
8 Paper Code
All the code for reproducing the experiments presented in this paper is available in the following two links. The first contains the forked version of Scikit-LearnPedregosa et al. [2011] which implements the era splitting models, and the second contains all the notebooks to recreate the figures and experiments from this research.